# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v2

@@ 3.1.1. Outcome Quality
Creativity. Creative tasks have the widest degree of output freedom. Representative crowdsourcing papers for creative objectives include Flash Teams (Retelny et al., 2014), VisiBlends (Chilton et al., 2019), and Mechanical Novel (Kim et al., 2017). These papers tackle creative writing (André et al., 2014; Kim et al., 2017; Bernstein et al., 2010; Yao et al., 2023; Yang et al., 2022; Mirowski et al., 2023; Zhang et al., 2023a), idea generation (Little et al., 2010a; Kim et al., 2023b; Wu et al., 2023c), and prototyping (Lee et al., 2017; Kulkarni et al., 2014; Valentine et al., 2017; Retelny et al., 2014; Park et al., 2022). Prior work finds that on creative tasks, workflows can help achieve higher quality (André et al., 2014; Kim et al., 2017; Yao et al., 2023; Yang et al., 2022; Mirowski et al., 2023; Kim et al., 2023b; Lee et al., 2017; Park et al., 2022) and less constrained outputs (Liu et al., 2018; Hahn et al., 2016; Lee et al., 2017; Drapeau et al., 2016; Wang et al., 2023b; Park et al., 2022). Workflows may enable novice crowdworkers (Pietrowicz et al., 2013) or users (Wang et al., 2023b; Lin et al., 2014a; André et al., 2013; Wu et al., 2023c; Valentine et al., 2017; Chilton et al., 2019; Kim and Monroy-Hernandez, 2016) to perform a creative task with an otherwise high barrier of entry. Important qualities for creative tasks include coherence (Yao et al., 2023; Yang et al., 2022; Kim et al., 2017) and aligning to the user’s creative vision (Salehi et al., 2017). On creative tasks, a workflow is often better suited for creating and lengthening the output, rather than proofreading (Kim et al., 2017; Huang et al., 2017; Little et al., 2010b; Mirowski et al., 2023), so these outputs are sometimes seen as a starting point for editing, rather than a finished product (Kim and Monroy-Hernandez, 2016; Park et al., 2022; Long et al., 2023; Mirowski et al., 2023; Nebeling et al., 2016).
3.1.2. Resource constraints
Workflow objectives are constrained by both monetary expense and latency requirements.
Expense. The expense of a workflow refers to paying crowdworkers or paying for an API call to an LLM. Workflows are often more expensive than simpler approaches (e.g., Chain of Thought prompting) (Zharikova et al., 2023; Yao et al., 2023; Wang et al., 2022; Du et al., 2023; Gero et al., 2023; Hahn et al., 2016; Huffaker et al., 2020; Lin et al., 2014a; Dai et al., 2010; Marcus et al., 2011a, b; Nair et al., 2023a). However, the cost is often less than hiring professionals (Ambati et al., 2012; Christoforaki and Ipeirotis, 2014; Zaidan and Callison-Burch, 2011; Liem et al., 2011; Fleming et al., 2023), pretraining models (Kim et al., 2023b; Levine et al., 2022; Yao et al., 2023; Huang et al., 2023b; Cobbe et al., 2021), or completing the task yourself (Zhang et al., 2022; Wang et al., 2023b).
Latency. Another common constraint is the time for the workflow to complete (Lasecki et al., 2015; Kulkarni et al., 2014; Marcus et al., 2011a, b). Many workflows have lower latency than performing the same task manually, since work can be parallelized across multiple workers (Teevan et al., 2016; Lasecki et al., 2015; Chilton et al., 2014; Guo et al., 2016; Lee et al., 2017; André et al., 2013; Pietrowicz et al., 2013; Lin et al., 2014a; Gouravajhala et al., 2018; Kulkarni et al., 2014).
3.2. Strategies in the design space
Several strategies can be applied to achieve workflow objectives. We categorize the strategies used in prior work into three groups: response diversity, quality control, and workflow flexibility (Figure 4).
3.2.1. Response diversity
Workflows can source multiple, diverse responses for the same subtask, which can promote either accuracy or creativity (Retelny et al., 2014; Kim et al., 2017; Nebeling et al., 2016; André et al., 2014; LaToza et al., 2014; Teevan et al., 2016). Crowdworkers have different viewpoints (Chen et al., 2019; Ambati et al., 2012; Lasecki et al., 2013) and sourcing perspectives from expert crowdworkers (Chilton et al., 2014) or from workers with various skill levels (Mahyar et al., 2018) can extend this variety. For LLMs, leveraging their stochastic nature (e.g., with a higher temperature parameter setting) or using different models or prompt variants (including role-based prompts) can source a variety of responses (Park et al., 2022; Wu et al., 2023c; Arora et al., 2022). In some cases one can automatically calculate the optimal mix of LLMs for a task (Madaan et al., 2023a).
Diversity for accuracy. When aggregated or compared, diverse responses can improve accuracy and robustness to low quality outputs, outliers, and spammers (Liem et al., 2011; Du et al., 2023; Zaidan and Callison-Burch, 2011; Huang et al., 2018; Lasecki et al., 2013; Huffaker et al., 2020; Reppert et al., 2023; Ambati et al., 2012; Hahn et al., 2016).
Three plots, each of which has several charts, each chart with a Y axis of Prevalence and two bars, one labeled Crowd and one labeled LLM. Plot 1: A set of two bar charts side by side, entitled Response Diversity. The first chart is labeled Diversity for Accuracy, with Crowd at 63.2% and LLM at 41.0%. The second chart is labeled Diversity for Creativity, with Crowd at 44.1% and LLM at 17.9%. Plot 2: A set of five bar charts side by side, entitled Quality Control. The first chart is labeled Appropriate Subtasks, with Crowd at 25.0% and LLM at 15.4%. The second chart is labeled Worker Instruction, with Crowd at 33.8% and LLM at 69.2%. The third chart is labeled Verification Step, with Crowd at 39.7% and LLM at 61.5%. The fourth chart is labeled Disagreement Resolution, with Crowd at 66.2% and LLM at 28.2%. The fifth chart is labeled User Revision, with Crowd at 29.4% and LLM at 12.8%. Plot 3: A set of four bar charts side by side, entitled Workflow Flexibility. The first chart is labeled Task Flexible, with Crowd at 27.9% and LLM at 48.7%. The second chart is labeled Task Specific, with Crowd at 72.1% and LLM at 51.3%. The third chart is labeled Dynamic, with Crowd at 51.5% and LLM at 51.3%. The fourth chart is labeled Rigid, with Crowd at 48.5% and LLM at 48.7%.
Diversity for creativity. Sourcing diverse responses encourages divergent and diverse thinking, helping improve the quality of creative outputs and to rapidly ideate (Liang et al., 2023; Wu et al., 2023a; Luther et al., 2015; Lasecki et al., 2015; Huang et al., 2017; Nebeling et al., 2016; Teevan et al., 2016). Furthermore, different workers can fill in each other’s deficits to create a more rounded final product (Kim and Monroy-Hernandez, 2016; Zhang et al., 2012).
3.2.2. Quality control
Errors can occur from the worker’s capabilities (Wu et al., 2022b; Yang et al., 2022; Xie et al., 2023) or can be induced by the workflow itself. Workflows can suffer from cascading errors (Bernstein et al., 2011; Wu et al., 2022a; Kittur et al., 2011; Goto et al., 2016; Little et al., 2010b; Bernstein et al., 2010; Little et al., 2010a; Wu et al., 2022b; Du et al., 2023) or produce incoherent outputs (Yang et al., 2022; Kim et al., 2017; Hahn et al., 2016; Mirowski et al., 2023; Wu et al., 2022a; Hahn et al., 2016). Aggregations may also induce disfluencies (Kim and Monroy-Hernandez, 2016; Bernstein et al., 2010; Hahn et al., 2016) or omit intended parts of a task (Yang et al., 2022). While all elements of a workflow should support quality outputs, some support quality control more directly: appropriate subtasks, worker instructions, verification steps, disagreement resolution mechanisms, and user revision.
Appropriate subtasks. Compared to entire tasks, subtasks are often easier to complete, verify, and correct (Ambati et al., 2012; Schick et al., 2022; Wu et al., 2022b; Yang et al., 2022; Kim et al., 2023b; Creswell et al., 2022; Creswell and Shanahan, 2022). Simplifying tasks can increase quality (Chilton et al., 2013; Wu et al., 2022b; Lasecki et al., 2013; Pietrowicz et al., 2013). Another strategy is to adapt the subtask to fit the worker’s capabilities (Chang et al., 2017; Mohanty et al., 2019; Pietrowicz et al., 2013; Arora et al., 2022). For example, researchers found that crowdworkers were better at generating predictive features than at estimating if a feature is predictive; so they adapted the workflow accordingly (Cheng and Bernstein, 2015).
Worker instructions. Properly instructing workers and identifying “good” workers can improve quality (Bragg et al., 2018; Liu et al., 2016). Both crowdworkers and LLMs benefit from task examples, namely screening and tutorial tasks for crowdworkers (Teevan et al., 2016; Lasecki et al., 2015; Chen et al., 2019; Lasecki et al., 2013; Pietrowicz et al., 2013; Drapeau et al., 2016; Bernstein et al., 2010) and few-shot examples for LLMs (Nair et al., 2023a; Gero et al., 2023). A key difficulty is communicating the user’s intent through language. Crowdworkers tend to need less explicitly defined instruction than LLMs, but can still struggle with vague or ambiguous instructions (Wu et al., 2023c). For both crowdworkers and LLMs, small changes in instructions or prompts can lead to dramatically different outcomes, impacting the accuracy of the final result (Wu et al., 2022b; Kittur et al., 2012; McDonnell et al., 2016; Wu et al., 2022a; Mirowski et al., 2023; Arora et al., 2022; Huang et al., 2023c; Wu et al., 2023c; Parameswaran et al., 2023; Gero et al., 2023). Additionally, different LLMs may respond differently to the same input prompt (Parameswaran et al., 2023).
Communicating context to workers can help with improving the quality of the output (Luther et al., 2015; Wu et al., 2022a; Teevan et al., 2016; Chilton et al., 2019; Hahn et al., 2016; Ambati et al., 2012; Kulkarni et al., 2012; Bernstein et al., 2010). The crowdsourcing literature contains many forms of contextual communication, from input text (Ambati et al., 2012; Kulkarni et al., 2012; Bernstein et al., 2010), to workflow goal tracking (Mirowski et al., 2023; Yang et al., 2022; Kim et al., 2017), to explicit communication between users and workers (Salehi et al., 2017). For LLMs, context may be included as part of the prompt, although LLMs have the additional barrier of an input token limit (Park et al., 2022; Mirowski et al., 2023).
Verification steps. Workflows can incorporate their own troubleshooting by explicitly dedicating a step for verification (Wu et al., 2023a, c; Zharikova et al., 2023; Kulkarni et al., 2012; Ambati et al., 2012; Schick et al., 2022). Others instead make all production editable at all steps so workers can fix errors in a flexible manner (Liu et al., 2018; LaToza et al., 2014). Automated methods of flagging malicious workers, gold-standard validation steps, or requiring a certain agreement threshold among workers can also provide verification (Lasecki et al., 2015; Parameswaran et al., 2023; Huffaker et al., 2020).
Disagreement resolution. Resolving disagreements among diverse responses can improve the quality of the aggregate outcome. In parallel and iterative workflows, errors from prior passes can be found and fixed or rejected (Zaidan and Callison-Burch, 2011; Hahn et al., 2016) and differing opinions can be merged (Kittur et al., 2011; Luther et al., 2015). Debate-based workflows can promote reflection on disagreements (Kobayashi et al., 2018; Chen et al., 2019; Du et al., 2023; Liang et al., 2023). Using different models and roles in debate shows promise (Du et al., 2023), as an LLM judge in a debate was found to more likely agree with its same underlying model (Liang et al., 2023).
User revision. Incorporating the user (the requester, in crowdsourcing terms) into some subtasks can also improve quality. Users can dynamically answer questions and provide feedback (Salehi et al., 2017) or monitor dynamic workflows to catch cascading errors (Kulkarni et al., 2012). Users can be called upon when necessary with a “panic button” that workers can invoke (Kulkarni et al., 2012), or by automatically surfacing ties to a user (Pietrowicz et al., 2013). With a shared interface between users and all workers, users can intervene when needed (Nebeling et al., 2016; Lasecki et al., 2015). Users may review at intermediate steps to edit or regenerate sections of the workflow (Park et al., 2022; Kim et al., 2023b; Zhang et al., 2023a). A workflow may also include multiple iterations with user revisions between each iteration (Willett et al., 2012; Cheng and Bernstein, 2015; Kulkarni et al., 2014; Park et al., 2022; Kim et al., 2023b). Users can also directly manipulate the final output to improve quality. This approach surfaces uncertainty in data as a feature for the user to explore, rather than hiding it in aggregation (Chang et al., 2017; Willett et al., 2012; Cheng and Bernstein, 2015). They can sort and filter the returned data (Willett et al., 2012; Cheng and Bernstein, 2015; Mohanty et al., 2019), use a slider to control output characteristics like text length (Bernstein et al., 2010), and consider category labels at multiple levels of precision (Chang et al., 2017).
3.2.3. Workflow flexibility
Workflows may or may not be task-specific, with either rigid or dynamic architectures.
@@ 3.2.3. Workflow flexibility
Task specificity. Some workflows are highly specialized to the specifications of the task (e.g., visual blends (Wang et al., 2023b; Chilton et al., 2019), masking faces in images (Alshaibani et al., 2020; Kaur et al., 2017), and programming (Huang et al., 2023b; LaToza et al., 2014)). Others generalize to a variety of tasks (Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Kulkarni et al., 2012; Wu et al., 2023a). For example, AutoGen (Wu et al., 2023a) uses the same conversation-based workflow for coding as well as chess games, math problems, and retrieval-based question answering. Task-specific workflows support fewer outcome types but can incorporate more tailored support of specific outcome qualities.
Architecture rigidity. Rigid workflows do not change their architecture depending on the input, while dynamic workflows generate parts of the workflow architecture at runtime (Valentine et al., 2017; Kim and Monroy-Hernandez, 2016; Kulkarni et al., 2014; Chen et al., 2019; Kittur et al., 2011; Huang et al., 2023c, b). Changes in dynamic workflows may be made by the crowd (LaToza et al., 2014; Valentine et al., 2017; Guo et al., 2016; Retelny et al., 2014), user (Retelny et al., 2014; Nebeling et al., 2016; Valentine et al., 2017; Kulkarni et al., 2012), or an algorithm (Bragg et al., 2013; Chen et al., 2019; Dai et al., 2013; Goto et al., 2016). Generally, architecture rigidity supports objectives that require more transparency and consistency, while architecture flexibility especially supports quality outcomes when tasks are difficult to predefine.
@@ 3.3. Tactics in the design space
While strategies help reach workflow objectives, more granular tactics are needed to implement them. Tactics need to consider the workflow architecture, the subtasks they contain, and the choice of workers for those subtasks (Figure 5).
Three plots, each of which has several charts, each chart with a Y axis of Prevalence and two bars, one labeled Crowd and one labeled LLM. Plot 1: A set of five bar charts side by side, entitled Architecture. The first chart is labeled Sequential, with Crowd at 100% and LLM at 100%. The second chart is labeled Branching, with Crowd at 10.3% and LLM at 12.8%. The third chart is labeled Dynamic, with Crowd at 52.9% and LLM at 51.3%. The fourth chart is labeled Communicative, with Crowd at 39.7% and LLM at 17.9%. The fifth chart is labeled Redundant, with Crowd at 85.3% and LLM at 64.1%. Plot 2: A set of five bar charts side by side, entitled Subtasks. The first chart is labeled Generate, with Crowd at 91.2% and LLM at 97.4%. The second chart is labeled Evaluate, with Crowd at 79.4% and LLM at 61.5%. The third chart is labeled Transform, with Crowd at 64.7% and LLM at 48.7%. The fourth chart is labeled Focus, with Crowd at 17.6% and LLM at 28.2%. The fifth chart is labeled Partition, with Crowd at 14.7% and LLM at 23.1%. Plot 3: A set of three bar charts side by side, entitled Workers. The first chart is labeled Crowdworkers, with Crowd at 100% and LLM at 0%. The second chart is labeled LLMs, with Crowd at 0% and LLM at 100%. The third chart is labeled User, with Crowd at 30.9% and LLM at 10.3%.

### v3

@@ 4.1.2. Accuracy vs creativity
Creativity. Creative tasks have the widest degree of output freedom. Representative crowdsourcing papers for creative objectives include Flash Teams (Retelny et al., 2014), VisiBlends (Chilton et al., 2019), and Mechanical Novel (Kim et al., 2017). These papers tackle creative writing (André et al., 2014; Kim et al., 2017; Bernstein et al., 2010; Yao et al., 2023; Yang et al., 2022; Mirowski et al., 2023; Zhang et al., 2023a), idea generation (Little et al., 2010a; Kim et al., 2023b; Wu et al., 2023c), and prototyping (Lee et al., 2017; Kulkarni et al., 2014; Valentine et al., 2017; Retelny et al., 2014; Park et al., 2022). Prior work finds that on creative tasks, workflows can help achieve higher quality (André et al., 2014; Kim et al., 2017; Yao et al., 2023; Yang et al., 2022; Mirowski et al., 2023; Kim et al., 2023b; Lee et al., 2017; Park et al., 2022) and less constrained outputs (Liu et al., 2018; Hahn et al., 2016; Lee et al., 2017; Drapeau et al., 2016; Wang et al., 2023; Park et al., 2022). Workflows may enable novice crowdworkers (Pietrowicz et al., 2013) or users (Wang et al., 2023; Lin et al., 2014a; André et al., 2013; Wu et al., 2023c; Valentine et al., 2017; Chilton et al., 2019; Kim and Monroy-Hernandez, 2016) to perform a creative task with an otherwise high barrier of entry. Important qualities for creative tasks include coherence (Yao et al., 2023; Yang et al., 2022; Kim et al., 2017) and aligning to the user’s creative vision (Salehi et al., 2017). On creative tasks, a workflow is often better suited for creating and lengthening the output, rather than proofreading (Kim et al., 2017; Huang et al., 2017; Little et al., 2010b; Mirowski et al., 2023), so these outputs are sometimes seen as a starting point for editing, rather than a finished product (Kim and Monroy-Hernandez, 2016; Park et al., 2022; Long et al., 2023; Mirowski et al., 2023; Nebeling et al., 2016).
Both accuracy and creativity. Many tasks require some degree of accuracy and some degree of creativity. In some tasks, there is not one correct answer, but there are restrictions on what outputs are correct. Crowdsourcing papers for these tasks include Cascade (Chilton et al., 2013), Flock (Cheng and Bernstein, 2015), Revolt (Chang et al., 2017), and Storia (Kim and Monroy-Hernandez, 2016). Example tasks include categorizing data (categories must be accurate, but there are multiple options and unclear edge cases) (Bragg et al., 2013; Chilton et al., 2013; André et al., 2013; Chilton et al., 2014; Cheng and Bernstein, 2015; Chang et al., 2017), summarizing text (summaries must stay faithful, but have degrees of freedom in how to write the summary) (Kim and Monroy-Hernandez, 2016; Pavel et al., 2014; Bernstein et al., 2010; Nair et al., 2023a, b), and coding (programs must work, but there are multiple approaches to problem solving) (LaToza et al., 2014; Kulkarni et al., 2014; Wu et al., 2023a; Huang et al., 2023b, a). The goals of workflows for these tasks are to align with expert outputs (Kim et al., 2014; Liu et al., 2018; Chilton et al., 2013; Salehi et al., 2017; Park et al., 2022; Kim et al., 2023b; Fleming et al., 2023) or to enable users to choose among multiple correct possible output options. Using workflows for these tasks can result in an output that is human-understandable (Cheng and Bernstein, 2015), that better represents what would be important for people to know (Salisbury et al., 2017), and that surfaces the uncertainty among the different possible correct options (Chang et al., 2017).
4.1.3. Task specificity vs generality
@@
Some workflows are highly specialized to the specifications of the task (e.g., visual blends (Wang et al., 2023; Chilton et al., 2019), masking faces in images (Alshaibani et al., 2020; Kaur et al., 2017), and programming (Huang et al., 2023b; LaToza et al., 2014)). Others generalize to a variety of tasks (Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Kulkarni et al., 2012; Wu et al., 2023a). For example, AutoGen (Wu et al., 2023a) uses the same conversation-based workflow for coding as well as chess games, math problems, and retrieval-based question answering. Task-specific workflows support fewer outcome types but can incorporate more tailored support than workflows that generalize to many tasks.
Title: Tactics, the building blocks of workflows. One row each for workers, subtasks, and architectures, with the category and (a short description). The workers row contains: Crowdworkers (people on an online platform), LLMs (generative large language models), user (the person deploying the workflow). The subtasks row contains: generate (create content), evaluate (judge content), transform (update content), focus (select content) and partition (split tasks). The architecture row contains: sequential (call LLMs in sequence), branching (diverge on paths), redundant (multiple workers repeat subtask), dynamic (Initialize parts of the workflow at runtime), and communicative (facilitate interaction among workers) Each subtask category is connected with a generic example instruction. Generate: ”Write a caption”. Evaluate: ”Is this answer grammatical?”. Transform: ”Improve the given idea”. Focus: ”Select regions for future work:. Partition: ”Divide the task into subtasks”.
@@ 4.2. Tactics in the design space
Designers use workflow tactics to implement workflows. Tactics need to consider the workers, the subtasks the workers complete, and the connections of subtasks into workflow architectures (Figure 3).
4.2.1. Workers
Workers may be crowdworkers, LLMs, or the user.
Crowdworkers. These workers are usually recruited from crowdworking platforms like Amazon Mechanical Turk or UpWork (formerly known as oDesk). Platforms allow for requests of worker qualifications, although these credentials are sometimes hard to validate (Ahmad et al., 2011; Zaidan and Callison-Burch, 2011; Valentine et al., 2017; Kulkarni et al., 2014). Methods like retainer models can reduce the latency of accessing crowdworkers on these platforms (Bernstein et al., 2011). Workers can also be recruited from the public, including community members (Mahyar et al., 2018), conference committees (Chilton et al., 2014; André et al., 2013), and online deaf communities (Bragg et al., 2021).
Crowdworkers have limitations that overlap with LLM limitations. Prior work suggests that approximately 30% of outcomes from Amazon Mechanical Turk are low quality (Bernstein et al., 2010), and may include spam or off-topic responses (Butler et al., 2017; Lasecki et al., 2013; Alshaibani et al., 2020; Huang et al., 2018). Crowdworkers also struggle with following long and complex instructions, as people can overly focus on the beginning or end of a given text (Hahn et al., 2016; Chilton et al., 2013; Lasecki et al., 2015; Chen et al., 2019). Crowdworkers can be biased; for example, in one study, workers performed significantly better in identifying photos of their own race compared to other races (Mohanty et al., 2019).
LLMs. There are many possible LLMs of many sizes and expenses (Bursztyn et al., 2022). Models have differing strengths and respond differently to prompts (Mirowski et al., 2023; Wu et al., 2023c; Parameswaran et al., 2023; Liang et al., 2023). LLM outputs can be adjusted through parameter changes (e.g., temperature) (Park et al., 2022; Parameswaran et al., 2023), fine-tuning (Creswell et al., 2022; Creswell and Shanahan, 2022; Cobbe et al., 2021; Li et al., 2023; Bursztyn et al., 2022), and expressing roles through prompts (Park et al., 2022; Du et al., 2023).
Some LLM limitations overlap with those of crowdworkers. LLMs can provide erroneous outputs: they are known to confidently output incorrect answers (Parameswaran et al., 2023; Wang et al., 2023; Du et al., 2023) or hallucinations (Nair et al., 2023a; Mirowski et al., 2023; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023) and for struggling with self-consistency and logical reasoning (Creswell et al., 2022; Creswell and Shanahan, 2022; Yao et al., 2023; Parameswaran et al., 2023; Liang et al., 2023; Xie et al., 2023). LLMs can ignore large fractions of long prompts or focus mostly on recent parts of the given context (Parameswaran et al., 2023; Du et al., 2023; Wu et al., 2023c; Madaan et al., 2023b). LLMs also output biased results, a widespread concern (Park et al., 2022; Wu et al., 2023a; Huang et al., 2023a; Long et al., 2023). For example, work on screenplay generation found stereotypical outputs, with multiple instances of gender biases (Mirowski et al., 2023). Furthermore, LLM training on human data increases the risk of accidental plagiarism (Mirowski et al., 2023).
LLMs also have unique limitations. On creative tasks, LLMs sometimes produce outputs that are perceived as robotic (Long et al., 2023) and that can lack the understanding, nuance, and subtext that people can provide (Mirowski et al., 2023). Furthermore, models can get stuck in generative loops, degenerating text quality (Parameswaran et al., 2023). LLMs also hallucinate more than crowds (Nair et al., 2023a; Mirowski et al., 2023; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023) and may struggle to change their outputs through self-reflection (Liang et al., 2023).
User. Finally, users of the workflow can also be involved to revise or guide the process. Users can be novices using the workflow to adapt to a steep learning curve (Kulkarni et al., 2014; Pietrowicz et al., 2013; Lin et al., 2014a; Chilton et al., 2019; Wang et al., 2023), or experts using the workflow as a support tool (Mirowski et al., 2023; Nebeling et al., 2016; Mohanty et al., 2019). Involving users in completing subtasks grants them greater control at the expense of higher effort (Stretcu et al., 2023). However, even with user involvement, prior work finds mixed results on how well tools retain the user’s sense of agency and ownership (Kim et al., 2023b; Mirowski et al., 2023).
4.2.2. Subtasks
Generate. Most workflows involve at least one step in which workers generate content. This generation may be writing words or code (Pavel et al., 2014; Weir et al., 2015; Zhang et al., 2022; Li et al., 2023; Zhang et al., 2023a) brainstorming and ideation (Chilton et al., 2019; Lasecki et al., 2013; Mahyar et al., 2018; Huang et al., 2018; Long et al., 2023; Yao et al., 2023), or answering a question or solving a problem (Kobayashi et al., 2018; Wu et al., 2023b; Xie et al., 2023; Eisenstein et al., 2022). Generation tasks may also ask for labels or annotations on given content (Guo et al., 2017, 2019; Chilton et al., 2014; Chang et al., 2017; Bragg et al., 2013), to translate between languages (Ambati et al., 2012; Zaidan and Callison-Burch, 2011), to contribute to a conversation with other workers (Mahyar et al., 2018; Wang et al., 2022; Liang et al., 2023; Du et al., 2023), to demonstrate a behavior (Bragg et al., 2021; Lee et al., 2017), or to source content from external sources (Kim et al., 2014; Chilton et al., 2019; Hahn et al., 2016; Kim and Monroy-Hernandez, 2016; Pietrowicz et al., 2013; Huang et al., 2015; Luther et al., 2015; Christoforaki and Ipeirotis, 2014). Some subtasks also require an explanation along with the generated content (Chang et al., 2017; Kim et al., 2017; Cheng and Bernstein, 2015; Drapeau et al., 2016; Schick et al., 2022; Gero et al., 2023; Fleming et al., 2023).
Evaluate. One of the most common evaluation tasks is choosing the best among multiple possible options (Guo et al., 2017; Weir et al., 2015; Guo et al., 2019; Cheng and Bernstein, 2015; Kittur et al., 2012; Liu et al., 2018). Other subtasks may include rating and ranking among multiple options (Pavel et al., 2014; Salehi et al., 2017; Marcus et al., 2011a; Huffaker et al., 2020; Willett et al., 2012; Mahyar et al., 2018) or a pairwise comparison (Mohanty et al., 2019; Little et al., 2010b, a; Marcus et al., 2011a, b; Parameswaran et al., 2023; Bursztyn et al., 2022). Evaluation subtasks may also determine if another round of generation needs to occur (Wu et al., 2022b; LaToza et al., 2014; Wang et al., 2022; Liang et al., 2023).
Workflows can decide what moves forward through majority vote (Huang et al., 2017; Dai et al., 2010; Lasecki et al., 2015; Liu et al., 2018; Kim et al., 2014; Guo et al., 2019), other aggregations of voting tasks (e.g., expectation maximization (Drapeau et al., 2016; Parameswaran et al., 2023), weighted mean (Mohanty et al., 2019), weak supervision (Arora et al., 2022)), returning the highest ranked or rated items (Pavel et al., 2014; Huffaker et al., 2020), or automated methods. These automated methods include measuring agreement from the generation step (Liu et al., 2018; Hahn et al., 2016; Huffaker et al., 2020; Chang et al., 2017; Huang et al., 2015; Nair et al., 2023b; Bernstein et al., 2011) and using non-AI checks like semantic embeddings and rule-based heuristics (Parameswaran et al., 2023; Wang et al., 2023; Ambati et al., 2012). There are also subtasks that seek critiques and reviews from the user (Nebeling et al., 2016; Schick et al., 2022) or another worker (Kim et al., 2017; Kulkarni et al., 2014; Valentine et al., 2017; Christoforaki and Ipeirotis, 2014; Mahyar et al., 2018; Schick et al., 2022; Parameswaran et al., 2023; Madaan et al., 2023b; Liang et al., 2023). These subtasks may ask for feedback or verify if errors or other qualities exist (Huang et al., 2023b; Bragg et al., 2013; Christoforaki and Ipeirotis, 2014; Chilton et al., 2019; Weir et al., 2015; Huang et al., 2015; Kim et al., 2014; Gero et al., 2023; Huang et al., 2017; Bernstein et al., 2010; Yang et al., 2022; Wu et al., 2023c; Kittur et al., 2011; Liang et al., 2023; Kulkarni et al., 2012; André et al., 2013; Butler et al., 2017).
Transform. Other subtasks transform given data by merging or improving it. Some subtasks merge multiple streams of incoming content. This merge could be algorithmic (Alshaibani et al., 2020; Ahmad et al., 2011; Alshaibani and Quinn, 2021; Teevan et al., 2016; Chilton et al., 2019; Kittur et al., 2011) or accomplished by the worker (Kittur et al., 2011; Wu et al., 2023c, 2022b; Lee et al., 2017; LaToza et al., 2014; Ambati et al., 2012; Teevan et al., 2016; Kittur et al., 2012; Wu et al., 2023c; Pietrowicz et al., 2013; Luther et al., 2015; Kulkarni et al., 2012). Other subtasks improve content generated by previous steps, sometimes by incorporating feedback from evaluation steps. Improving on previously generated content can come in many forms: updating the format of the content (Huang et al., 2023b; Wu et al., 2022b; Eisenstein et al., 2022; Wang et al., 2022; Arora et al., 2022), pruning bad results (Liu et al., 2018; Gero et al., 2023), adding missing results (Huang et al., 2023b; Liu et al., 2018; Weir et al., 2015), correcting errors (Huang et al., 2023b; Yang et al., 2022; Drapeau et al., 2016; Kobayashi et al., 2018; Gouravajhala et al., 2018), and performing direct improvements (Madaan et al., 2023b; Hahn et al., 2016; Mahyar et al., 2018; Lee et al., 2017; Huffaker et al., 2020; Nebeling et al., 2016; Christoforaki and Ipeirotis, 2014; Little et al., 2010b; Liem et al., 2011; Dai et al., 2010; Schick et al., 2022; Kim et al., 2017; Bernstein et al., 2010; Guo et al., 2019; McDonnell et al., 2016).
Focus. One set of subtasks looks to select the attention of the upcoming tasks within a broader context. For example, this step may find patches of the input that need editing, extract information from context, or focus on common keywords (Bernstein et al., 2011; Huffaker et al., 2020; Gero et al., 2023; Creswell and Shanahan, 2022; Creswell et al., 2022; Nair et al., 2023b; Kim et al., 2017).
Partition. Some subtasks transform a task by breaking it up into smaller pieces (Kittur et al., 2011; LaToza et al., 2014; Kulkarni et al., 2012; Wu et al., 2023c; Zhou et al., 2022b; Valentine et al., 2017; Wu et al., 2022b; Huang et al., 2023c). For example, the Divide step in the Turkomatic workflow asks workers to divide a given subtask into simpler subtasks. The workflow calls this step recursively until the subtask complexity is valued (by a separate crowdworker) at a specified cost (Kulkarni et al., 2012).
Three columns headed by ”If…”, ”Then…” and ”With…”. The ”If” column contains different possible challenges relating to achieving objectives. The ”Then” column lists all of the strategies. There are arrows connecting the ”If” column to the ”Then” column, with arrows from ”navigating a large creative space” objectives challenge to the adaptive architectures, diverse responses, and user guidance strategies. The ”With” column lists all of the tactics. There are arrows connecting the ”Then” column to the ”With” column. There are many arrows, and they correspond to the tactics used for strategies as defined in the text and other figures.